# Matching-Based Policy Learning

### v1

@@ 2.2 Matching-Based Advantage Function
Our proposed approach is inspired by an alternative expression of the advantage function:
$A(\pi)=E\left[\left(2\pi(X)-1\right)\left(Y(1)-Y(0)\right)\right]=E\left[\left(2\pi(X)-1\right)\operatorname{sign}\left(Y(1)-Y(0)\right)\left|Y(1)-Y(0)\right|\right].$
The second equality indicates that, in order to maximize $A(\pi)$, one should assign each unit to the treatment group in which its potential outcome can be larger; meanwhile, a more significant discrepancy between the pair of potential outcomes contributes to a greater influence in the treatment allocation. Building on this intuition, we seek to develop an estimator of $A(\pi)$ based on the pair of potential outcomes. However, since only one of the pair of potential outcomes is observable, we utilize matching estimator to impute the missing counterfactual outcome.
We leverage the matching estimator proposed by Abadie and Imbens (2006), where they employed matching with replacement, allowing each unit to be used as a match multiple times. Specifically, for any $x\in\mathcal{X}$ and some positive definite matrix $V$, let $\|x\|=(x^{\top}Vx)^{1/2}$ denote a vector norm. We define the matched set $\mathcal{J}_{M}(i)$ as the set that contains the indices of the first $M$ matches for unit $i$:
@@
Here, we ignore the ties in matching, which happens with probability zero if $X$ is continuously distributed (Abadie and Imbens, 2006). In this paper, we consider the Mahalanobis metric $V_{\text{maha}}$, which is commonly used in practice:
$V_{\text{maha}}=\left(\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\bar{X})(X_{i}-\bar{X})^{\top}\right)^{-1},\text{ where }\bar{X}=\frac{1}{n}\sum_{i=1}^{n}X_{i}.$
The Mahalanobis metric is not the only choice, see Abadie and Imbens (2006) for discussion on other metrics. The matching estimator of potential outcomes is given by
$\hat{Y}_{i}(0)=\left\{\begin{aligned} &Y_{i}&\text{ if }W_{i}=0,\\ &\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}Y_{j}&\text{ if }W_{i}=1,\end{aligned}\right.\quad\text{ and }\quad\hat{Y}_{i}(1)=\left\{\begin{aligned} &\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}Y_{j}&\text{ if }W_{i}=0,\\ &Y_{i}&\text{ if }W_{i}=1.\end{aligned}\right.$
For each unit $i$, the pair of matching estimators $\left(\hat{Y}_{i}(1),\hat{Y}_{i}(0)\right)$ mimics the pair of potential outcomes $\left(Y_{i}(1),Y_{i}(0)\right)$. Finally, we propose the matching-based advantage function:
$\hat{A}_{match}(\pi)=\frac{1}{n}\sum^{n}_{i=1}\left(2\pi(X_{i})-1\right)\left(\hat{Y}_{i}(1)-\hat{Y}_{i}(0)\right).$ (1)
In the following, we will establish the consistency and asymptotic normality of the matching-based estimator $\hat{A}_{match}(\pi)$, with subsequent discussion of its limitation and the proposal for improvement. Let $M_{i}=(M_{i1},\ldots,M_{in})^{\top}$ denote the $n$-dimensional vector indicating the units in $\mathcal{J}_{M}(i)$, where $M_{ij}=1$ if unit $j\in\mathcal{J}_{M}(i)$, and $M_{ij}=0$ otherwise, i.e., $M_{ij}=I\left\{j\in\mathcal{J}_{M}(i)\right\}$. For convenience of further analysis, we formulate $\hat{A}_{match}(\pi)$ as a linear combination of $\{Y_{i}\}_{i=1}^{n}$ after basic manipulations.
@@
$\hat{A}_{match}(\pi)=\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left[(2\pi(X_{i})-1)+\frac{K_{M}(\pi,i)}{M}\right]Y_{i},$
where $K_{M}(\pi,i)=\sum^{n}_{j=1}\left(2\pi(X_{j})-1\right)M_{ji}=\sum_{j:M_{ji}=1}\left(2\pi(X_{j})-1\right)$.
It is noteworthy that $\{K_{M}(\pi,i)\}_{i=1}^{n}$ is the key component for asymptotic analysis. The properties of $\hat{A}_{match}(\pi)$ resemble those of the matching estimator for $\tau$ proposed by Abadie and Imbens (2006):
$\hat{\tau}=\frac{1}{n}\sum^{n}_{i=1}\left(\hat{Y}_{i}(1)-\hat{Y}_{i}(0)\right)=\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left(1+\frac{K_{M}(i)}{M}\right)Y_{i},$
where $K_{M}(i)=\sum_{j=1}^{n}M_{ji}$ is the number of times unit $i$ has been matched. When $\pi(x)\equiv 1$, it follows that $\hat{A}_{match}(\pi)=\hat{\tau}$ and $K_{M}(\pi,i)=K_{M}(i)$. Therefore, the matching estimator $\hat{\tau}$ is a special case of our proposed estimator. Furthermore, we proceed to show that $K_{M}(\pi,i)$ has the same order as $K_{M}(i)$ under the following assumption.
@@
The $p$-dimensional vector $X$ is continuously distributed on a convex and compact support $\mathcal{X}_{0}\subset\mathbb{R}^{p}$, and its density is bounded and bound away from zero on $\mathcal{X}_{0}$.
Assumption 1 is a standard regularity condition for the matching estimator of average causal effect (Abadie and Imbens, 2006), supposing $X$ is continuously distributed. In contrast, discrete covariates with finite support points can be addressed by analyzing the estimation of the advantage function within subsamples. However, for high-dimensional and continuous $X$, the assumption that its density is bound away from zero may be too stringent (Otsu and Rai, 2017).
@@
Under Assumption 1, for any $\pi\in\Pi$, $K_{M}(\pi,i)=O_{p}(1)$ and $E[K_{M}(\pi,i)^{q}]$ is bounded uniformly in $n$ for any $q>0$.
Analyzing the asymptotic properties of $\hat{A}_{match}(\pi)$ may be challenging due to the dependence among $\{K_{M}(\pi,i)\}^{n}_{i=1}$. Inspired by Abadie and Imbens (2006), we decompose $\hat{A}_{match}(\pi)$ into several components for further analysis.
@@
$\bar{A}(\pi)=\frac{1}{n}\sum^{n}_{i=1}(2\pi(X_{i})-1)(\mu(X_{i},1)-\mu(X_{i},0)),$
$E_{M}(\pi)=\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left[(2\pi(X_{i})-1)+\frac{K_{M}(\pi,i)}{M}\right]\epsilon_{i},$
$B_{M}(\pi)=\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)(2\pi(X_{i})-1)\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}(\mu(X_{i},1-W_{i})-\mu(X_{j},1-W_{i})).$ (2)
Let $\mathbf{X}$ denote the $n\times p$ matrix whose $i$th row is $X^{\top}_{i}$, and $\mathbf{W}$ is the $n\times 1$ vector with the $i$th element equal to $W_{i}$. It follows that $B_{M}(\pi)=E\left[\hat{A}_{match}(\pi)-\bar{A}(\pi)\mid\mathbf{X},\mathbf{W}\right]$ after some algebra, referred to as the conditional bias relative to $\bar{A}(\pi)$. Utilizing the decomposition in Lemma 3, we proceed to derive the large sample properties of $\hat{A}_{match}(\pi)$ by analyzing its components separately. First, we derive the order of $B_{M}(\pi)$ in the following lemma.
@@
Lemma 4 demonstrates that, although the conditional bias is asymptotically negligible, its convergence rate can be slow with a moderately large $p$. Let $V^{\tau(X)}(\pi)=E\left[(2\pi(X)-1)\tau(X)-A(\pi)\right]^{2}$. The conditional variance of $E_{M}(\pi)$ given $\mathbf{X}$ and $\mathbf{W}$ is
$\text{Var}\left[E_{M}(\pi)\mid\mathbf{X},\mathbf{W}\right]=\frac{1}{n^{2}}\sum^{n}_{i=1}\left[(2\pi(X_{i})-1)+\frac{K_{M}(\pi,i)}{M}\right]^{2}\sigma^{2}(X_{i},W_{i}),$
and the corresponding normalized conditional variance is denoted by $V^{E}(\pi)=n\text{Var}\left[E_{M}(\pi)\mid\mathbf{X},\mathbf{W}\right]$. Under the following regularity conditions, we establish the consistency and asymptotic normality of $\hat{A}_{match}(\pi)$.
@@
(ii) For $w=0,1$, E$\left[Y^{4}\mid X=x,W=w\right]$ is bounded uniformly in $\mathcal{X}_{0}$.
(iii) For any $\pi\in\Pi$, $V^{E}(\pi)$ bounded away from zero in the Cartesian product space $\mathcal{X}_{0}^{n}\times\{0,1\}^{n}$.
Assumption 2(i) is a mild condition for restricting the smoothness and boundness of conditional mean and variance. Assumption 2(ii) is a moment condition for the central limit theorem. Assumption 2(iii) is analogous to the Assumption 4(iii) in Abadie and Imbens (2006), which guarantees the nonsingularity of $V^{E}(\pi)$ uniformly in $\Pi$.
@@
(i) Under Assumptions 1 and 2(i), then for any $\pi\in\Pi$,
$\hat{A}_{match}(\pi)-A(\pi)\stackrel{{\scriptstyle p}}{{\rightarrow}}0.$
(ii) Under Assumptions 1 and 2, then for any $\pi\in\Pi$,
$\left(V^{\tau(X)}(\pi)+V^{E}(\pi)\right)^{-1/2}\sqrt{n}\left(\hat{A}_{match}(\pi)-B_{M}(\pi)-A(\pi)\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N(0,1).$
Theorem 1 shows that the matching-based estimator $\hat{A}_{match}(\pi)$ converges in probability to the advantage function $A(\pi)$ regardless of covariates dimension $p$. In addition, $\hat{A}_{match}(\pi)$ is $\sqrt{n}$-asymptotically normal distributed after subtracting the conditional bias $B_{M}(\pi)$. Furthermore, leveraging Lemma 4 and Theorem 1, we establish that
@@
When the covariate is scalar, $\hat{A}_{match}(\pi)$ possesses $\sqrt{n}$-asymptotic normality. However, the convergence rate can be slow with a moderately large $p$, where $B_{M}(\pi)$ becomes the leading term. Therefore, it is imperative to improve the estimator $\hat{A}_{match}(\pi)$ that eliminates the conditional bias and thus remedies the slow rate.
2.3 Bias-Corrected Matching-Based Advantage Function
To alleviate the conditional bias in $\hat{\tau}$, Abadie and Imbens (2011) utilized a power series estimator of the outcome mean model $\mu(x,w)$ to develop a bias-corrected matching estimator. Specifically, let $\lambda=\left(\lambda_{1},\ldots,\lambda_{p}\right)^{\top}$ denote a $p$-dimensional vector of nonnegative integers with the norm $\|\lambda\|_{1}=\sum_{j=1}^{p}\lambda_{j}$. Define that $x^{\lambda}=\prod_{j=1}^{p}(x_{j})^{\lambda_{j}}$, where $x_{j}$ is the $j$th element of $p$-dimensional vector $x$. For the series $\{\lambda(j)\}_{j=1}^{\infty}$ containing all distinct vectors with nondecreasing $\|\lambda(j)\|_{1}$, define the $K$-dimensional vector $p^{K}(x)=\left(x^{\lambda(1)},\ldots,x^{\lambda(K)}\right)^{\top}$. The power series estimator (Newey, 1997) of $\mu(x,w)$ is given by
$\hat{\mu}(x,w)=\left(p^{K}(x)\right)^{\top}\left(\underset{i:W_{i}=w}{\sum}p^{K}(X_{i})\left(p^{K}(X_{i})\right)^{\top}\right)^{-}\underset{i:W_{i}=w}{\sum}p^{K}(X_{i})Y_{i},$
where $(\cdot)^{-}$ is a generalized inverse. Finally, the bias-corrected matching estimator of potential outcomes is defined as
$\tilde{Y}_{i}(0)=\left\{\begin{aligned} &Y_{i}&\text{ if }W_{i}=0,\\ &\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}\left(Y_{j}+\hat{\mu}(X_{i},0)-\hat{\mu}(X_{j},0)\right)&\text{ if }W_{i}=1,\end{aligned}\right.$

### v2

@@ 2.2 Matching-Based Advantage Function
We rewrite the advantage function in the following expression:
$A(\pi)$ $=$ $E\left[\left(2\pi(X)-1\right)\left\{Y(1)-Y(0)\right\}\right]$
$=$ $E\left[\left\{2\pi(X)-1\right\}\operatorname{sign}\left\{Y(1)-Y(0)\right\}\left|Y(1)-Y(0)\right|\right].$
The second equality indicates that, in order to maximize $A(\pi)$ one should assign each unit to the treatment group in which its potential outcome is larger. Meanwhile, a more significant discrepancy between the pair of potential outcomes contributes to a greater influence in the treatment allocation. Building on this observation, we seek to develop an estimator of $A(\pi)$ based on the pair of potential outcomes. Half of the potential outcomes are missing, and we utilize matching techniques to impute the missing outcomes.
We leverage the matching estimator in Abadie and Imbens (2006), where NN matching with replacement is in use. Specifically, for any $x\in\mathcal{X}$ and some positive definite matrix $V$, let $\|x\|=(x^{\top}Vx)^{1/2}$ denote a vector norm. We define the matched set $\mathcal{J}_{M}(i)$ as the set that contains the indices of the first $M$ matches for unit $i$:
@@
Throughout, we assume there are no ties in matching, which is satisfied if $X$ is continuously distributed (Abadie and Imbens, 2006). Tackling ties will be theoretically more demanding and it is an interesting direction of future work. In this paper, we consider the commonly used Mahalanobis metric: $\|x\|=(x^{\top}V_{\text{maha}}x)^{1/2}$, where $V_{\text{maha}}=\left\{n^{-1}\sum_{i=1}^{n}(X_{i}-\bar{X})(X_{i}-\bar{X})^{\top}\right\}^{-1}$ and $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$. See Abadie and Imbens (2006) for discussion on other metrics. The matching estimator of potential outcomes is given by
$\hat{Y}_{i}(0)=\left\{\begin{aligned} &Y_{i},&\text{ if }W_{i}=0,\\ &\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}Y_{j},&\text{ if }W_{i}=1,\end{aligned}\right.\quad\text{ and }\quad\hat{Y}_{i}(1)=\left\{\begin{aligned} &\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}Y_{j},&\text{ if }W_{i}=0,\\ &Y_{i},&\text{ if }W_{i}=1.\end{aligned}\right.$
For each unit $i$, the pair of matching estimators $\{\hat{Y}_{i}(1),\hat{Y}_{i}(0)\}$ mimics the pair of potential outcomes $\{Y_{i}(1),Y_{i}(0)\}$. Finally, we propose the matching-based advantage function:
$\hat{A}_{match}(\pi)=\frac{1}{n}\sum^{n}_{i=1}\left\{2\pi(X_{i})-1\right\}\left\{\hat{Y}_{i}(1)-\hat{Y}_{i}(0)\right\}.$ (1)
In the following, we establish the consistency and asymptotic normality of the matching-based estimator $\hat{A}_{match}(\pi)$, followed by discussion of its limitation and the proposal of improvement. Let $M_{i}=(M_{i1},\ldots,M_{in})^{\top}$ denote the $n$-dimensional vector indicating the units in $\mathcal{J}_{M}(i)$, where $M_{ij}=I\left\{j\in\mathcal{J}_{M}(i)\right\}$. That is, $M_{ij}=1$ if unit $j\in\mathcal{J}_{M}(i)$, and $M_{ij}=0$ otherwise. For convenience of further analysis, we rewrite $\hat{A}_{match}(\pi)$ as a linear combination of $\{Y_{i}\}_{i=1}^{n}$ using elementary algebra.
@@
$\hat{A}_{match}(\pi)=\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left[\{2\pi(X_{i})-1\}+\frac{K_{M}(\pi,i)}{M}\right]Y_{i},$
where $K_{M}(\pi,i)=\sum^{n}_{j=1}\{2\pi(X_{j})-1\}M_{ji}=\sum_{j:M_{ji}=1}{2\pi(X_{j})-1}$.
It is noteworthy that $\{K_{M}(\pi,i)\}_{i=1}^{n}$ is the key component for theoretical developments. Recall that the matching estimator for the ATE $\tau$ in Abadie and Imbens (2006) is
$\hat{\tau}=\frac{1}{n}\sum^{n}_{i=1}\left\{\hat{Y}_{i}(1)-\hat{Y}_{i}(0)\right\}=\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left\{1+\frac{K_{M}(i)}{M}\right\}Y_{i},$
where $K_{M}(i)=\sum_{j=1}^{n}M_{ji}$ is the number of times unit $i$ has been matched. The ATE estimator $\hat{\tau}$ and the advantage function estimator $\hat{A}_{match}(\pi)$ have similar expressions. In particular, when $\pi(x)\equiv 1$, it follows that $\hat{A}_{match}(\pi)=\hat{\tau}$ and $K_{M}(\pi,i)=K_{M}(i)$. Therefore, the matching estimator $\hat{\tau}$ is a special case of our proposed estimator with the “treat-all” policy. In the following, we show that the properties of $\hat{A}_{match}(\pi)$ resemble those of $\hat{\tau}$.
@@
The $p$-dimensional vector $X$ is continuously distributed on a convex and compact support $\mathcal{X}_{0}\subset\mathbb{R}^{p}$. Its density is upper bounded. Moreover, its density is bounded away from zero on $\mathcal{X}_{0}$.
Assumption 1 is a standard regularity condition for the matching estimator of average causal effect (Abadie and Imbens, 2006). Discrete covariates with finite supports can be addressed by analyzing the estimation of the advantage function within subsamples. For high-dimensional and continuous $X$, the assumption that its density is bounded away from zero may be too stringent (Otsu and Rai, 2017). Future work on relaxing this assumption is warranted.
@@
Under Assumption 1, for any $\pi\in\Pi$, $K_{M}(\pi,i)=O_{p}(1)$ and $E[K_{M}(\pi,i)^{q}]$ is bounded uniformly for any $n$ and $q>0$.
This lemma says that $K_{M}(\pi,i)$ has the same order as $K_{M}(i)$ under Assumption 1. Analyzing the asymptotic properties of $\hat{A}_{match}(\pi)$ is challenging due to the dependence among the terms $\{K_{M}(\pi,i)\}^{n}_{i=1}$. To alleviate the difficulty, we decompose $\hat{A}_{match}(\pi)$ into several components for further analysis as in Abadie and Imbens (2006).
@@
$\bar{A}(\pi)$ $=$ $\frac{1}{n}\sum^{n}_{i=1}\{2\pi(X_{i})-1\}\{\mu(X_{i},1)-\mu(X_{i},0)\},$
$E_{M}(\pi)$ $=$ $\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left[\{2\pi(X_{i})-1\}+\frac{K_{M}(\pi,i)}{M}\right]\epsilon_{i},$
$B_{M}(\pi)$ $=$ $\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\{2\pi(X_{i})-1\}\left[\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}\{\mu(X_{i},1-W_{i})-\mu(X_{j},1-W_{i})\}\right].$ (2)
Let $\mathbf{X}$ be an $n\times p$ matrix where the $i$th row is $X^{\top}_{i}$. Let $\mathbf{W}$ be the $n\times 1$ vector with the $i$th element equal to $W_{i}$. It follows that $B_{M}(\pi)=E[\hat{A}_{match}(\pi)-\bar{A}(\pi)\mid\mathbf{X},\mathbf{W}]$ after some algebra, referred to as the conditional bias term relative to $\bar{A}(\pi)$. Utilizing the decomposition in Lemma 3, we proceed to derive the large sample properties of $\hat{A}_{match}(\pi)$ by analyzing its three components respectively. First, we derive the order of $B_{M}(\pi)$ in the following lemma.
@@
Lemma 4 suggests that the conditional bias $B_{M}(\pi)$ is asymptotically negligible, but the convergence rate is quite slow when the covariate dimension $p$ is moderate or large. The conditional variance of $E_{M}(\pi)$ given $\mathbf{X}$ and $\mathbf{W}$ is:
$\text{Var}\left[E_{M}(\pi)\mid\mathbf{X},\mathbf{W}\right]=\frac{1}{n^{2}}\sum^{n}_{i=1}\left[\{2\pi(X_{i})-1\}+\frac{K_{M}(\pi,i)}{M}\right]^{2}\sigma^{2}(X_{i},W_{i}).$
Define $V^{E}(\pi)=n\text{Var}\left[E_{M}(\pi)\mid\mathbf{X},\mathbf{W}\right]$.
@@
(ii) For $w=0,1$, E$\left[Y^{4}\mid X=x,W=w\right]$ is upper bounded uniformly in $\mathcal{X}_{0}$.
(iii) For any $\pi\in\Pi$, $V^{E}(\pi)$ is bounded away from zero for any $(\mathbf{X},\mathbf{W})\in\mathcal{X}_{0}^{n}\times\{0,1\}^{n}$.
Assumption 2(i) is a mild condition on the smoothness and boundedness of the conditional mean and the conditional variance. Assumption 2(ii) is a moment condition for the central limit theorem. Assumption 2(iii) is analogous to the Assumption 4(iii) in Abadie and Imbens (2006), which guarantees the non-singularity of $V^{E}(\pi)$ uniformly in $\Pi$.
Let $V^{\tau(X)}(\pi)=E[\{2\pi(X)-1\}\tau(X)-A(\pi)]^{2}$. We establish the asymptotic properties for $\hat{A}_{match}(\pi)$ in the following theorem.
@@
(i) Under Assumptions 1 and 2(i), then for any $\pi\in\Pi$, as $n\rightarrow\infty$,
$\hat{A}_{match}(\pi)-A(\pi)\stackrel{{\scriptstyle p}}{{\longrightarrow}}0.$
(ii) Under Assumptions 1 and 2, then for any $\pi\in\Pi$, as $n\rightarrow\infty$,
$\left\{V^{\tau(X)}(\pi)+V^{E}(\pi)\right\}^{-1/2}\sqrt{n}\left\{\hat{A}_{match}(\pi)-B_{M}(\pi)-A(\pi)\right\}\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(0,1).$
Theorem 1 shows that the matching-based estimator $\hat{A}_{match}(\pi)$ converges in probability to the advantage function $A(\pi)$ regardless of the covariate dimension $p$. Furthermore, $\hat{A}_{match}(\pi)$ is asymptotically normally distributed after subtracting the conditional bias term $B_{M}(\pi)$. Importantly, Theorem 1 together with Lemma 4 lead to the following rate of convergence result.
Corollary 1.
Under Assumptions 1 and 2, for any $\pi\in\Pi$,
@@
Therefore, when the covariate is scalar or bivariate, $\hat{A}_{match}(\pi)$ is $\sqrt{n}$-consistent. However, the convergence rate can be very slow with a moderately large $p$, because $B_{M}(\pi)$ becomes the leading term. Therefore, it is imperative to eliminate the conditional bias and thus remedy the slow rate for the matched-based estimator of the advantage function.
2.3 Bias-Corrected Matching Estimator of the Advantage Function
We utilize the idea of bias correction in Abadie and Imbens (2011) to improve the matching estimator of the advantage function. Let $\lambda=\left(\lambda_{1},\ldots,\lambda_{p}\right)^{\top}$ be a $p$-dimensional vector of nonnegative integers with the norm $\|\lambda\|_{1}=\sum_{j=1}^{p}\lambda_{j}$. Define that $x^{\lambda}=\prod_{j=1}^{p}(x_{j})^{\lambda_{j}}$, where $x_{j}$ is the $j$th element of $p$-dimensional vector $x$. For the series $\{\lambda(j)\}_{j=1}^{\infty}$ containing all distinct vectors with nondecreasing $\|\lambda(j)\|_{1}$, define the $K$-dimensional vector $p^{K}(x)=\left(x^{\lambda(1)},\ldots,x^{\lambda(K)}\right)^{\top}$. The power series estimator (Newey, 1997) of $\mu(x,w)$ is given by
$\hat{\mu}(x,w)=\left(p^{K}(x)\right)^{\top}\left\{\underset{i:W_{i}=w}{\sum}p^{K}(X_{i})\left(p^{K}(X_{i})\right)^{\top}\right\}^{-}\underset{i:W_{i}=w}{\sum}p^{K}(X_{i})Y_{i},$
where $(\cdot)^{-}$ is a generalized inverse. The bias-corrected matching estimator of the potential outcomes (Abadie and Imbens, 2011) is
$\tilde{Y}_{i}(0)=\left\{\begin{aligned} &Y_{i},&\text{ if }W_{i}=0,\\ &\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}\left\{Y_{j}+\hat{\mu}(X_{i},0)-\hat{\mu}(X_{j},0)\right\},&\text{ if }W_{i}=1,\end{aligned}\right.$